OpenStudy (goformit100):

A person walks along a circular path of radius 5.00 m. If the person walks around one half of the circle, find (a) the magnitude of the displacement vector and (b) how far the person walked. (c) What is the magnitude of the displacement if the person walks all the way around the circle?

terenzreignz (terenzreignz):

Conceptual questions... if you walk halfway around the circle, you end up on the other side, right?

terenzreignz (terenzreignz):

 Now walk halfway around the circle...

OpenStudy (goformit100):

ok

terenzreignz (terenzreignz):

 IF you go halfway, you'd have walked half the circumference of the circle... \[\huge \frac12 C = \frac12 \cdot 2\pi r= \pi r = 5\pi\]